# Categorical Time-Reversal Symmetries

### 30 Mar 2026 v1

@@ I Introduction
The modern understanding of symmetries in quantum many-body systems and quantum field theory has expanded well beyond the traditional setting of groups. This has been initiated by the deep insight of [1], that topological defects are symmetries. One of the key physically relevant implications of these extensions is that symmetries need not have inverses and form groups, but can form more general, non-invertible structures. These are subsumed under the name of categorical symmetries and were recently reviewed in [2, 3, 4, 5].
@@
The main purpose of this paper is to explore the categorical generalization, including identifying a mathematical home for such anti-unitary categorical symmetries and studying their implications in the context of the categorical classification of phases and phase transitions [10, 11, 12, 13, 14]. We show that the natural setting for categorical anti-unitary symmetries are real fusion categories. The mathematical literature on this is surprisingly scarce [15, 16], and we will be providing some of the background on these categories in the following.
@@
with $\mathcal{C}_{1}$ the linear sector and $\mathcal{C}_{T}$ the anti-linear. Crucially, we will argue that for a quantum system, to have a complex Hilbert space as the state space, it is the Galois-real fusion categories that have an interpretation as symmetry categories – $\mathbb{R}$-real categories, essentially have $\mathbb{R}$ acting on the local state space, which we disregard as a consistent setup for quantum theories. The two ${\mathbb{Z}}_{2}^{T}$-graded components of a Galois-real fusion category naturally correspond to the linear and anti-linear sectors of a generalized anti-linear symmetry. In particular, $\mathcal{C}_{1}$ alone is an ordinary fusion category over $\mathbb{C}$}.
@@
Gapped Phases. We propose a module category approach for the classification of gapped phases with generalized anti-linear symmetries. This is akin to the unitary case studied in [10]. Gapped phases with symmetry are in 1-1 correspondence with module categories over the now Galois-real symmetry category, which characterize how the symmetry category acts on the underlying gapped phase. E.g. we can reproduce with this approach the known twisted group-cohomology classification for invertible $G^{T}$-symmetries. The module category approach also provides an explicit description of the categories of defects within each gapped phase and between different phases in terms of module functors. We compute the categorical structure of all gapped phases with group-like symmetry ${\mathbb{Z}}_{4}^{T}$ – see Figure 1 – and $S_{3}^{T}={\mathbb{Z}}_{3}\rtimes{\mathbb{Z}}_{2}^{T}$ – see Table 1 – including the categories of defects within each phase and the categories of domain walls between phases. As a consequence of the analysis, we establish the known Morita equivalence  [^2] It is known that $\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{T}$ with mixed anomaly is related to $\mathbb{Z}_{4}^{T}$ by gauging [19]. This relation is usually established through field-theoretic arguments or explicit lattice constructions. To the best of our knowledge, however, a proof in terms of Morita equivalence has not appeared prior to the present work.:
@@
The SPT phases with time-reversal symmetry have a long history, going back to the Haldane chain. One of the challenges studying these phases is the absence of a local or string order parameter, for a discussion of this see [20, 21, 7, 22, 23, 24]. The ${\mathbb{Z}}_{2}^{T}$ phase can be detected by computing a partition function on real projective space $\mathbb{R}P^{2}$, and on a space with boundary the edge modes can be identified with a Kramers doublet. However there seems to be no direct bulk order parameter. Other (1+1)d gapped phases with time-reversal symmetry and the interplay of this with gapped phase classification has been studied in [25, 26, 27].
Dual Symmetries. One of the advantage of the categorical formulation of symmetries and phases is that it allows for direct and rigorous computation of dualities (generalized gauging) between symmetries. We extend the theory of generalized gauging [28, 29] to include anti-linear symmetries, and use the Morita theory for real fusion categories to define dual anti-linear symmetries. As an application of our formalism. We establish the following family of dualities/gauge-relations generated by gauging central subgroups.
@@
${\mathbb{Z}}_{2}^{T}$-Enriched SymTFT. The SymTFT [30, 31, 32, 33] has been a key tool in carrying out the Categorical Landau Paradigm Program, for gapped phases in 1+1d [12, 34, 35, 36, 37, 38, 39] and 2+1d [40, 41, 42, 43, 44], including second order, symmetric, phase transitions [45, 14, 46, 47, 48, 13, 49, 50, 51], and mixed phases [39, 52, 53], and most recently continuous internal symmetries [54, 55, 56, 57, 58]. The fundamental idea is to start with a theory with a global symmetry $\mathcal{C}$, coupling this to background fields, and gauge in one dimension higher. This protocol applies to both internal symmetries as well as continuous spacetime symmetries [58]. However, for discrete spacetime symmetries such as time-reversal and translations, it was argued in [19], that instead of gauging these symmetries, the SymTFT for internal symmetries is merely enriched by the spacetime symmetries, in the sense of [59, 60]. From an unoriented Turaev-Viro theory perspective some of these aspects appeared in [61].
@@
Plan of the Paper. In Section II we motivate real fusion categories directly from simple (1+1)(1+1)d systems with time-reversal symmetry. Using a ${\mathbb{Z}}_{2}^{T}$-symmetric spin chain, we derive the two basic categories associated with time-reversal, namely $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})\simeq\mathsf{Vec}_{\mathbb{R}}$ and $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$, from the categories of symmetric defects in the trivial and symmetry-broken phases. We then revisit the Haldane chain, explain the appearance of the quaternion algebra $\mathbb{H}$ and the real category $\mathsf{Vec}_{\mathbb{H}}$, discuss the lack of string order parameters for anti-linear symmetries, and formulate the module-category description of (1+1)(1+1)d ${\mathbb{Z}}_{2}^{T}$-phases. In Section III we discuss the general theory of real fusion categories. We distinguish the two types, $\mathbb{R}$-real and Galois-real fusion categories, explain their different physical roles, and review the theory of module categories and Morita equivalence over real fusion categories. In Section Section IV we illustrate these structures with concrete examples. In particular, we discuss the group-theoretical categories associated with anti-linear groups $G^{T}$, the categories $\mathsf{Rep}(G^{T})$, non-invertible time-reversal symmetries of Tambara–Yamagami type, and the semi-direct product categories $\mathcal{C}\rtimes\mathbb{Z}_{2}^{T}$. In Section V we turn to the classification of (1+1)(1+1)d gapped phases with generalized anti-linear symmetries. We propose that such phases are classified by module categories over the Galois-real symmetry category, develop the corresponding notion of duality and gauging, and revisit the basic $\mathbb{Z}_{2}^{T}$ examples from this perspective. We then analyze the general structure of phases with group-like anti-linear symmetry and study in detail the examples of $\mathbb{Z}_{4}^{T}$ and $S_{3}^{T}=\mathbb{Z}_{3}\rtimes\mathbb{Z}_{2}^{T}$, including their defects, domain walls, and Morita dual symmetry categories. We conclude the section by proving the family of Motira equivalences generated by gauging finite unitary subgroups.
@@
We change gear in Section VI and we develop the $\mathbb{Z}_{2}^{T}$-enriched (or simply T-enriched) SymTFT framework. We explain the distinction between T-enrichment and gauging, formulate the bulk theory in terms of T-enriched topological order, and identify the relevant bulk defects with the Drinfeld center of a Galois-real fusion category. In Section VII we apply this framework to SymTFT quiches for anti-linear symmetries. We explain how a Galois-real symmetry category arises on a boundary of a $\mathbb{Z}_{2}^{T}$-enriched SymTFT, and show through examples how Morita equivalent anti-linear symmetries appear as different symmetry boundaries of the same bulk. The appendices collect some technical material, including additional discussion of the semi-direct product categories $\mathcal{C}\rtimes\mathbb{Z}_{2}^{T}$ and the proof of the gauging theorem. We conclude with some open questions in Section VIII.
@@ II Motivating Real Fusion Categories from ${\mathbb{Z}}_{2}^{T}$-symmetric Phases
One of the key points of this paper is that real fusion categories are the relevant mathematical structure underlying the study of systems with anti-linear symmetries. The goal of this section is to introduce and motivate from a physical picture, how real fusion categories naturally arise in the study of anti-linear symmetries, such as ${\mathbb{Z}}_{2}^{T}$.
The most basic example of an anti-linear symmetry is the time-reversal symmetry ${\mathbb{Z}}_{2}^{T}$. In this section we will derive two real fusion categories associated with it, denoted $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$ and $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})$, respectively, from a concrete 1 spin chain with ${\mathbb{Z}}_{2}^{T}$-symmetry. The notation is deliberately chosen to match with the $\mathsf{Vec}_{G},\penalty 10000\ \mathsf{Rep}(G)$ categories associated with a finite, internal, $\mathbb{C}$}-linear $G$-symmetry. The $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})$ category is in fact the same as $\mathsf{Vec}_{\mathbb{R}}$ the category of (finite dimensional) real vector spaces
@@
where $K$ means complex conjugation in the basis where $Z_{i}$ are diagonal and $U=\prod_{i}X_{i}$ is the usual ${\mathbb{Z}}_{2}$-symmetry generator in the transverse Ising model. The choice of the anti-unitary symmetry action is not unique and will not affect the conclusions in this section. With the choice of generator (II.3) which we will denote by ${\mathbb{Z}}_{2}^{T}$, the ground state is ${\mathbb{Z}}_{2}^{T}$-symmetric when $J<1$ and breaks ${\mathbb{Z}}_{2}^{T}$ spontaneously when $J>1$.
@@ II.1 ${\mathbb{Z}}_{2}^{T}$-Trivial Phase
The category of symmetric defects in a ${\mathbb{Z}}_{2}^{T}$-symmetric trivial phase is $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})\simeq\mathsf{Vec}_{\mathbb{R}}$. ∎
@@
We see that the space spanned by the operator $Z_{k}$ together with its ${\mathbb{Z}}_{2}^{T}$-action is equivalent to the complex plane with the standard complex conjugation action. In fact, any irreducible representation of ${\mathbb{Z}}_{2}^{T}$ is equivalent to the complex plane with ${\mathbb{Z}}_{2}^{T}$ acting as complex conjugation, $T$. This becomes the mathematical statement tha the category $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})$ of ${\mathbb{Z}}_{2}^{T}$-representations contains only one simple object, $({\mathbb{C}},T)$, up to equivalence.
@@
Therefore only multiplication by real numbers are allowed as maps of defects. We conclude that in the category of symmetric defects, we have $\text{End}(\mathbf{1})\simeq\mathbb{R}$. Put everything together we see that the category of symmetric defects in the trivial ${\mathbb{Z}}_{2}^{T}$-symmetric phase is a real fusion category generated by a single simple object $\mathbf{1}$ satisfying $\text{End}(\mathbf{1})\simeq\mathbb{R}$. This is nothing but the category $\mathsf{Vec}_{\mathbb{R}}$ of finite dimensional real vector spaces. In fact $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})$ is equivalent to $\mathsf{Vec}_{\mathbb{R}}$: a complex vector space with anti-unitary ${\mathbb{Z}}_{2}^{T}$-action is the same as a real vector space by taking the fixed points of the ${\mathbb{Z}}_{2}^{T}$-action:
@@ II.2 ${\mathbb{Z}}_{2}^{T}$ SSB-Phase
where $|\uparrow\rangle$ and $|\downarrow\rangle$ the eigenstates of $Z$ with eigenvalues $\pm 1$. The domain wall defect is created by the symmetric non-local operator
@@
To fully determine the structure of the category of symmetric defects we need to do more work. Next we consider the endomorphisms of the trivial defect $\Omega_{\mathbf{1}}$. Assume that there is a ${\mathbb{Z}}_{2}^{T}$-symmetric local operator $\mathcal{O}$ that acts on the ground state space as a general linear transformation
@@
One is eager to conclude that the category of symmetric defects is the (complex) fusion category $\mathsf{Vec}_{{\mathbb{Z}}_{2}}$, with the non-trivial simple object being the domain wall defect $\mathsf{D}$. However this is not the case and a subtle but crucial difference will be revealed by further analysis. Let us consider a domain wall $\mathsf{D}_{k}$ at site $k$, and a local symmetric operator $\mathcal{O}_{j}$ realizing the endomorphism $(z,w)\mapsto(\alpha z,\alpha^{*}w)$ at site $j$. One choice is $\mathcal{O}_{j}^{\alpha}=a\,\mathrm{id}+i\,b\,Z_{k},\penalty 10000\ \alpha=a+bi$. On the $|\uparrow\cdots\uparrow\rangle$ ground state, $\mathcal{O}_{j}$ acts as $\alpha\mathrm{id}_{\uparrow}$, while on the $|\downarrow\cdots\downarrow\rangle$ ground state it acts as $\alpha^{*}\mathrm{id}_{\downarrow}$. Let $j<k<l$, we notice the crucial relation
@@
which says that a local scalar $\alpha\in\text{End}(\mathbf{1})$ is complex conjugated when passes through the defect $\mathsf{D}$. This is expected: the defect $\mathsf{D}$ is a time-reversal domain wall, which should be the Wick-rotation of an anti-linear action, thus it should act on scalars as complex conjugation.
Another way to look at this is the following. The defect $\mathsf{D}_{k}$ is the same as $\mathbf{1}\otimes\mathsf{D}_{k}$, this means any endomorphism of the trivial defect $\alpha:\mathbf{1}\to\mathbf{1}$ gives rise to a endomorphism of the defect $\mathsf{D}_{k}$ by tensor product: $\alpha\mathrm{id}_{\mathbf{1}}\otimes\mathrm{id}_{\mathsf{D}}$. Concretely this means we obtain an endomorphism of $\mathsf{D}$ by putting the operator $\mathcal{O}_{j}$ to the left of the domain wall $\mathsf{D}_{k}$($j<k$). This leads to an endomorphism:
@@
Similarly, the defect $\mathsf{D}_{k}$ is the same as $\mathsf{D}_{k}\otimes\mathbf{1}$, therefore putting any endomorphism of $\mathbf{1}$ to the right of $\mathsf{D}_{k}$ also gives an endomorphism of $\mathsf{D}_{k}$. Concretely, acting with $\mathcal{O}_{j}$ on $\mathsf{D}_{k}$ with $i>k$ leads to

### 02 Apr 2026 v2

@@ I Introduction
The modern understanding of symmetries in quantum many-body systems and quantum field theory has expanded well beyond the traditional setting of groups. This was initiated by the insight of [1] that topological defects are symmetries. One of the key physically relevant implications of these extensions is that symmetries need not have inverses and form groups, but can form more general, non-invertible structures. These are subsumed under the name of categorical symmetries and were recently reviewed in [2, 3, 4, 5].
@@
The main purpose of this paper is to explore the categorical generalization, including identifying a mathematical home for such anti-unitary categorical symmetries and studying their implications in the context of the categorical classification of phases and phase transitions [10, 11, 12, 13, 14]. We show that the natural setting for categorical anti-unitary symmetries is real fusion categories. The mathematical literature on this is surprisingly scarce [15, 16], and we will be providing some of the background on these categories in the following.
@@
with $\mathcal{C}_{1}$ the linear sector and $\mathcal{C}_{T}$ the anti-linear. Crucially, we will argue that for a quantum system, to have a complex Hilbert space as the state space, it is the Galois-real fusion categories that have an interpretation as symmetry categories – $\mathbb{R}$-real categories essentially have $\mathbb{R}$ acting on the local state space, which we disregard as a consistent setup for quantum theories. The two ${\mathbb{Z}}_{2}^{T}$-graded components of a Galois-real fusion category naturally correspond to the linear and anti-linear sectors of a generalized anti-linear symmetry. In particular, $\mathcal{C}_{1}$ alone is an ordinary fusion category over $\mathbb{C}$}.
@@
Gapped Phases. We propose a module category approach for the classification of gapped phases with generalized anti-linear symmetries. This is akin to the unitary case studied in [10]. Gapped phases with symmetry are in one-to-one correspondence with module categories over the now Galois-real symmetry category, which characterize how the symmetry category acts on the underlying gapped phase. E.g. we can reproduce with this approach the known twisted group-cohomology classification for invertible $G^{T}$-symmetries. The module category approach also provides an explicit description of the categories of defects within each gapped phase and between different phases in terms of module functors. We compute the categorical structure of all gapped phases with group-like symmetry ${\mathbb{Z}}_{4}^{T}$ – see Figure 1 – and $S_{3}^{T}={\mathbb{Z}}_{3}\rtimes{\mathbb{Z}}_{2}^{T}$ – see Table 1 – including the categories of defects within each phase and the categories of domain walls between phases. As a consequence of the analysis, we establish the known Morita equivalence  [^2] It is known that $\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{T}$ with mixed anomaly is related to $\mathbb{Z}_{4}^{T}$ by gauging [19]. This relation is usually established through field-theoretic arguments or explicit lattice constructions. To the best of our knowledge, however, a proof in terms of Morita equivalence has not appeared prior to the present work.:
@@
The SPT phases with time-reversal symmetry have a long history, going back to the Haldane chain. One of the challenges studying these phases is the absence of a local or string order parameter, for a discussion, see [20, 21, 7, 22, 23, 24]. The ${\mathbb{Z}}_{2}^{T}$ phase can be detected by computing a partition function on real projective space $\mathbb{R}P^{2}$, and on a space with boundary the edge modes can be identified with a Kramers doublet. However, there seems to be no direct bulk order parameter. Other (1+1)d gapped phases with time-reversal symmetry and the interplay of this with gapped phase classification has been studied in [25, 26, 27].
Dual Symmetries. One of the advantages of the categorical formulation of symmetries and phases is that it allows for direct and rigorous computation of dualities (generalized gauging) between symmetries. We extend the theory of generalized gauging [28, 29] to include anti-linear symmetries, and use the Morita theory for real fusion categories to define dual anti-linear symmetries. As an application of our formalism, we establish the following family of dualities/gauge-relations generated by gauging central subgroups.
@@
${\mathbb{Z}}_{2}^{T}$-Enriched SymTFT. The SymTFT [30, 31, 32, 33] has been a key tool in carrying out the Categorical Landau Paradigm Program, for gapped phases in 1+1d [34, 12, 35, 36, 37, 38, 39, 40] and 2+1d [41, 42, 43, 44, 45, 46], including second order, symmetric, phase transitions [47, 14, 48, 49, 50, 13, 51, 52, 53], and mixed phases [40, 54, 55], and most recently continuous internal symmetries [56, 57, 58, 59, 60, 61]. The fundamental idea is to start with a theory with a global symmetry $\mathcal{C}$, coupling this to background fields, and gauge in one dimension higher. This protocol applies to both internal symmetries as well as continuous spacetime symmetries [60]. However, for discrete spacetime symmetries such as time-reversal and translations, it was argued in [19], that instead of gauging these symmetries, the SymTFT for internal symmetries is merely enriched by the spacetime symmetries, in the sense of [62, 63]. From an unoriented Turaev-Viro theory perspective some of these aspects appeared in [64].
@@
Plan of the Paper. In Section II we motivate real fusion categories directly from simple (1+1)(1+1)d systems with time-reversal symmetry. Using a ${\mathbb{Z}}_{2}^{T}$-symmetric spin chain, we derive the two basic categories associated with time-reversal, namely $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})\simeq\mathsf{Vec}_{\mathbb{R}}$ and $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$, from the categories of symmetric defects in the trivial and symmetry-broken phases. We then revisit the Haldane chain, explain the appearance of the quaternion algebra $\mathbb{H}$ and the real category $\mathsf{Vec}_{\mathbb{H}}$, discuss the lack of string order parameters for anti-linear symmetries, and formulate the module-category description of (1+1)(1+1)d ${\mathbb{Z}}_{2}^{T}$-phases. In Section III we discuss the general theory of real fusion categories. We distinguish the two types, $\mathbb{R}$-real and Galois-real fusion categories, explain their different physical roles, and review the theory of module categories and Morita equivalence over real fusion categories. In Section IV we illustrate these structures with concrete examples. In particular, we discuss the group-theoretical categories associated with anti-linear groups $G^{T}$, the categories $\mathsf{Rep}(G^{T})$, non-invertible time-reversal symmetries of Tambara–Yamagami type, and the semi-direct product categories $\mathcal{C}\rtimes\mathbb{Z}_{2}^{T}$. In Section V we turn to the classification of (1+1)(1+1)d gapped phases with generalized anti-linear symmetries. We propose that such phases are classified by module categories over the Galois-real symmetry category, develop the corresponding notion of duality and gauging, and revisit the basic $\mathbb{Z}_{2}^{T}$ examples from this perspective. We then analyze the general structure of phases with group-like anti-linear symmetry and study in detail the examples of $\mathbb{Z}_{4}^{T}$ and $S_{3}^{T}=\mathbb{Z}_{3}\rtimes\mathbb{Z}_{2}^{T}$, including their defects, domain walls, and Morita dual symmetry categories. We conclude the section by proving the family of Morita equivalences generated by gauging finite unitary subgroups.
@@
We change gears in Section VI and we develop the $\mathbb{Z}_{2}^{T}$-enriched (or simply T-enriched) SymTFT framework. We explain the distinction between T-enrichment and gauging, formulate the bulk theory in terms of T-enriched topological order, and identify the relevant bulk defects with the Drinfeld center of a Galois-real fusion category. In Section VII we apply this framework to SymTFT quiches for anti-linear symmetries. We explain how a Galois-real symmetry category arises on a boundary of a $\mathbb{Z}_{2}^{T}$-enriched SymTFT, and show through examples how Morita equivalent anti-linear symmetries appear as different symmetry boundaries of the same bulk. The appendices collect some technical material, including additional discussion of the semi-direct product categories $\mathcal{C}\rtimes\mathbb{Z}_{2}^{T}$ and the proof of the gauging theorem. We conclude with some open questions in Section VIII.
@@ II Motivating Real Fusion Categories from ${\mathbb{Z}}_{2}^{T}$-symmetric Phases
One of the key points of this paper is that real fusion categories are the relevant mathematical structure underlying the study of systems with anti-linear symmetries. The goal of this section is to explain, from a physical perspective, how real fusion categories naturally arise in the study of anti-linear symmetries, such as ${\mathbb{Z}}_{2}^{T}$.
The most basic example of an anti-linear symmetry is the time-reversal symmetry ${\mathbb{Z}}_{2}^{T}$. In this section we will derive two real fusion categories associated with it, denoted $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$ and $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})$, respectively, from a concrete 1D spin chain with ${\mathbb{Z}}_{2}^{T}$-symmetry. The notation is deliberately chosen to match with the $\mathsf{Vec}_{G},\penalty 10000\ \mathsf{Rep}(G)$ categories associated with a finite, internal, $\mathbb{C}$}-linear $G$-symmetry. The $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})$ category is in fact the same as $\mathsf{Vec}_{\mathbb{R}}$ the category of (finite dimensional) real vector spaces
@@
where $K$ means complex conjugation in the basis where $Z_{i}$ are diagonal and $U=\prod_{i}X_{i}$ is the usual ${\mathbb{Z}}_{2}$-symmetry generator in the transverse Ising model. The choice of the anti-unitary symmetry action is not unique and will not affect the conclusions in this section. With the choice of generator (II.3) which we will denote by ${\mathbb{Z}}_{2}^{T}$, the ground state is ${\mathbb{Z}}_{2}^{T}$-symmetric when $J<1/2$ and breaks ${\mathbb{Z}}_{2}^{T}$ spontaneously when $J>1/2$.
@@ II.1 ${\mathbb{Z}}_{2}^{T}$-Trivial Phase
The category of symmetric defects in a ${\mathbb{Z}}_{2}^{T}$-symmetric trivial phase is $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})\simeq\mathsf{Vec}_{\mathbb{R}}$.
@@
We see that the space spanned by the operator $Z_{k}$ together with its ${\mathbb{Z}}_{2}^{T}$-action is equivalent to the complex plane with the standard complex conjugation action. In fact, any irreducible representation of ${\mathbb{Z}}_{2}^{T}$ is equivalent to the complex plane with ${\mathbb{Z}}_{2}^{T}$ acting as complex conjugation, $T$. This becomes the mathematical statement that the category $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})$ of ${\mathbb{Z}}_{2}^{T}$-representations contains only one simple object, $({\mathbb{C}},T)$, up to equivalence.
@@
Therefore only multiplication by real numbers is allowed as maps of defects. We conclude that in the category of symmetric defects, we have $\text{End}(\mathbf{1})\simeq\mathbb{R}$. Put everything together we see that the category of symmetric defects in the trivial ${\mathbb{Z}}_{2}^{T}$-symmetric phase is a real fusion category generated by a single simple object $\mathbf{1}$ satisfying $\text{End}(\mathbf{1})\simeq\mathbb{R}$. This is nothing but the category $\mathsf{Vec}_{\mathbb{R}}$ of finite dimensional real vector spaces. In fact $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})$ is equivalent to $\mathsf{Vec}_{\mathbb{R}}$: a complex vector space with anti-unitary ${\mathbb{Z}}_{2}^{T}$-action is the same as a real vector space by taking the fixed points of the ${\mathbb{Z}}_{2}^{T}$-action:
@@ II.2 ${\mathbb{Z}}_{2}^{T}$ SSB-Phase
where $|\uparrow\rangle$ and $|\downarrow\rangle$ are the eigenstates of $Z$ with eigenvalues $\pm 1$. The domain wall defect is created by the symmetric non-local operator
@@
To fully determine the structure of the category of symmetric defects we need to do more work. Next we consider the endomorphisms of the trivial defect $\Omega$. Assume that there is a ${\mathbb{Z}}_{2}^{T}$-symmetric local operator $\mathcal{O}$ that acts on the ground state space as a general linear transformation
@@
One might conclude that the category of symmetric defects is the (complex) fusion category $\mathsf{Vec}_{{\mathbb{Z}}_{2}}$, with the non-trivial simple object being the domain wall defect $\mathsf{D}$. However, this is not the case and a subtle but crucial difference will be revealed by further analysis. Let us consider a domain wall $\mathsf{D}_{k}$ at site $k$, and a local symmetric operator $\mathcal{O}_{j}$ realizing the endomorphism $(z,w)\mapsto(\alpha z,\alpha^{*}w)$ at site $j$. One choice is $\mathcal{O}_{j}^{\alpha}=a\,\mathrm{id}+i\,b\,Z_{j},\penalty 10000\ \alpha=a+bi$. On the $|\uparrow\cdots\uparrow\rangle$ ground state, $\mathcal{O}_{j}$ acts as $\alpha\mathrm{id}_{\uparrow}$, while on the $|\downarrow\cdots\downarrow\rangle$ ground state it acts as $\alpha^{*}\mathrm{id}_{\downarrow}$. Let $j<k<l$, we notice the crucial relation
@@
which says that a local scalar $\alpha\in\text{End}(\mathbf{1})$ is complex conjugated when it passes through the defect $\mathsf{D}$. This is expected: the defect $\mathsf{D}$ is a time-reversal domain wall, which should be the Wick-rotation of an anti-linear action, thus it should act on scalars as complex conjugation.
Another way to look at this is the following. The defect $\mathsf{D}_{k}$ is the same as $\mathbf{1}\otimes\mathsf{D}_{k}$, this means any endomorphism of the trivial defect $\alpha:\mathbf{1}\to\mathbf{1}$ gives rise to an endomorphism of the defect $\mathsf{D}_{k}$ by tensor product: $\alpha\mathrm{id}_{\mathbf{1}}\otimes\mathrm{id}_{\mathsf{D}}$. Concretely this means we obtain an endomorphism of $\mathsf{D}$ by putting the operator $\mathcal{O}_{j}$ to the left of the domain wall $\mathsf{D}_{k}$($j<k$). This leads to an endomorphism:
@@
Similarly, the defect $\mathsf{D}_{k}$ is the same as $\mathsf{D}_{k}\otimes\mathbf{1}$, therefore putting any endomorphism of $\mathbf{1}$ to the right of $\mathsf{D}_{k}$ also gives an endomorphism of $\mathsf{D}_{k}$. Concretely, acting with $\mathcal{O}_{j}$ on $\mathsf{D}_{k}$ with $j>k$ leads to